VIBRATIONAL SPECTRAL STUDIES OF METAL(II) HALIDE COMPLEXES WITH 3-AMINOTOLUENE AND 4-AMINOTOLUENE

\underline{M. KUMRU}, A. ALTUN and K. GOLCUK, Fatih University, Faculty of Arts and Sciences, Department of Physics, 34900 Büyükcekmece, Istanbul, Turkey.